Fourdrops of a 0.45M imidazole solution were added to a solution in a cuvette containing ferrihemoglobin. The mass of the ferrihemoglobin solution in the cuvette was measured to be 2.170 g initially and 2.252 g after the addition of the stock imidazole solution. What is the molar concentration of imidazole in the final solution after the 4 drops of stock imidazole were added to the ferrihemoglobin solution in the cuvette?
Answer #1

Imidazole has an approximated density of 1.23 g/ml. However, it is hard to acknowledge the density of the ferrohemoglobin solution. We could assume it to be 1 g/ml, such as water, else we would have to obtain said information from the problem itself.

Knowing the density, we can calculate the volume of each solution, with the formula:

\rho = \frac{m}{V}, from which we can solve for Volume and get: V = \frac{m}{\rho}

As such, the volumes are:

Vimidazole = 2.252g - 2.170g -= 0.0667ml 1.23g/ml

2.2529 Vf-hemoglobin = = 2.252ml 1g/ml

Now we can use the following formula:

CiVi = CV, where C 1 and 2 are the concentrations of each solution and V 1 and 2 are the volumes of each solution. We can solve for C2 as:

Civi C2 = - (0.45M) 0.0667ml) 2 = 0.12M (2.252ml)

The concentration of imidazole is now about 0.12 M.